### v1

# QR-LoRA: Efficient and Disentangled Fine-tuning via QR Decomposition for Customized Generation

Abstract
Existing text-to-image models often rely on parameter fine-tuning techniques such as Low-Rank Adaptation (LoRA) to customize visual attributes. However, when combining multiple LoRA models for content-style fusion tasks, unstructured modifications of weight matrices often lead to undesired feature entanglement between content and style attributes. We propose QR-LoRA, a novel fine-tuning framework leveraging QR decomposition for structured parameter updates that effectively separate visual attributes. Our key insight is that the orthogonal Q matrix naturally minimizes interference between different visual features, while the upper triangular R matrix efficiently encodes attribute-specific transformations. Our approach fixes both Q and R matrices while only training an additional task-specific $\Delta R$ matrix. This structured design reduces trainable parameters to half of conventional LoRA methods and supports effective merging of multiple adaptations without cross-contamination due to the strong disentanglement properties between $\Delta R$ matrices. Experiments demonstrate that QR-LoRA achieves superior disentanglement in content-style fusion tasks, establishing a new paradigm for parameter-efficient, disentangled fine-tuning in generative models.
Figure 1: Given a single content and a single style image, we present QR-LoRA, a method to achieve efficient and disentangled control of content and style features through orthogonal decomposition. Our approach reduces trainable parameters while maintaining superior disentanglement properties, enabling flexible manipulation of visual attributes with enhanced initialization. Best viewed with zoom-in.
1 Introduction
Feature disentanglement and control in generative models have been a long-standing challenge, particularly in the domain of image synthesis where precise manipulation of visual attributes is crucial for practical applications. Recent advances in diffusion models [16, 53, 42, 47, 54, 40, 21, 73, 34, 35, 72] have demonstrated remarkable capabilities in generating high-quality images. These models [73, 72, 61, 18, 44, 25, 6, 48, 52, 15, 20, 19] have shown impressive progress in both content generation and style manipulation, enabling diverse applications from photorealistic image synthesis to artistic style transfer. The rapid development of foundation models [16, 54, 53, 21, 68] and adaptation techniques [17, 48, 38, 73, 72] have made customizing generative models for specific visual tasks increasingly accessible, marking a milestone in AI-driven creative applications.
Contemporary approaches to feature disentanglement can be broadly categorized into content-specific [48, 73, 38], style-specific [52, 15, 61], and joint content-style methods [49, 8, 31]. While single-task approaches have shown success in their respective domains, the increasing demand for simultaneous control over both content and style has led to the development of more comprehensive solutions. Among various model adaptation techniques, Low-Rank Adaptation (LoRA) [17] has emerged as a particularly promising approach for its parameter efficiency and versatility. Yet, current joint content-style methods based on LoRA face significant challenges in achieving effective feature disentanglement while maintaining parameter efficiency. Recent solutions like ZipLoRA [49] attempt to achieve effective disentanglement by learning orthogonal coefficients between independently trained LoRAs, highlighting the importance of orthogonality in feature separation. However, such approaches, along with methods like B-LoRA [8], ComposLoRA [74], CMLoRA [76], remain heavily dependent on model-specific architectures [47, 44] and domain constraints. While the pursuit of orthogonality provides valuable insights for feature disentanglement, the field still lacks a unified, model-agnostic approach that can directly learn orthogonal adaptation modules without relying on post-training coefficient optimization.
In this paper, we introduce QR-LoRA, a novel fine-tuning framework that leverages QR decomposition [11] for disentangled content and style control. Our approach is motivated by both theoretical foundations and empirical observations. Through empirical analysis in Figure 2, we observe that $Q$ matrices exhibit remarkably high similarity (approaching $1.0$) across different adaptation tasks, while $R$ matrices show relatively lower similarity. This stability of $Q$ matrices is theoretically supported by the minimal Frobenius norm property of orthogonal parameterization (detailed in Appendix), which ensures that the orthogonal basis vectors maintain consistent orthogonality while minimizing redundant transformations in the feature space. Based on these insights, we propose a novel initialization paradigm that maintains $Q$ as fixed orthogonal basis vectors spanning the linear transformation space, while introducing a $\Delta R$ mechanism to learn task-specific features, ensuring effective feature disentanglement while minimizing interference with the model’s original capabilities.
Our framework begins with singular value decomposition (SVD)-based [4, 10, 13, 50, 36, 29] core information extraction, followed by orthogonal basis construction through QR decomposition. The key innovation lies in our $\Delta R$ mechanism, which learns compact, task-specific transformations in the orthogonal space while preserving the model’s original capabilities. This approach not only achieves more precise feature disentanglement but also significantly reduces the number of trainable parameters. Unlike existing methods that rely on model-specific architectural analysis [8] or complex merging strategies for independently trained LoRAs [49], our method provides an elegant, model-agnostic solution that maintains consistent performance across various architectures. Extensive experiments demonstrate that QR-LoRA achieves superior disentanglement capabilities and faster convergence compared to conventional approaches, while requiring only half of the trainable parameters. Our main contributions are as follows:
  • A novel SVD-QR based initialization paradigm for LoRA that constructs orthogonal basis vectors through matrix decomposition, reducing trainable parameters by half while maintaining model performance.
  • A residual mechanism ($\Delta R$) that learns task-specific transformations in a fixed orthogonal space, enabling precise control with disentanglement properties between visual attributes for effective merging.
  • We demonstrate superior performance of QR-LoRA in content-style fusion tasks, with consistent results across various model architectures, highlighting its significant potential for broader applications.
2 Related Work
Parameter-Efficient Fine-tuning (PEFT) has emerged as a pivotal research area in large-scale generative models [64, 28, 39], particularly in adapting foundation models with minimal computational overhead. Various PEFT strategies have been proposed, including Adapter modules [55] that insert trainable layers, Prompt-tuning [27] that optimizes continuous prompts, and LoRA [17] that decomposes weight updates into low-rank matrices (i.e., $AB$). Among these, LoRA has demonstrated remarkable success in large language models, inspiring innovative designs such as HydraLoRA [57] with asymmetric structures that leverage shared $A$ matrices for global feature integration and task-specific $B$ matrices for capturing intrinsic components, effectively minimizing interference between different tasks. Recent advances like PISSA [36] leverage SVD-based initialization for accelerated convergence, while SVDiff [13] introduces a compact parameter space by fine-tuning singular values of weight matrices, significantly reducing model size while maintaining performance. SVFT [29] further advances the field by efficiently updating weights through sparse combinations of singular vectors, achieving remarkable performance with minimal trainable parameters. These advances in parameter-efficient learning have provided valuable insights for text-to-image model adaptation. However, the direct application of these techniques to visual generation tasks faces unique challenges due to the complex nature of visual feature disentanglement, especially in scenarios requiring fine-grained control over both content and style attributes.
Image customization in text-to-image models has evolved along several distinct technical trajectories. Model adaptation approaches [52, 72, 73, 38, 49, 31, 28] encompass various strategies from parameter-efficient modules to feature injection mechanisms and lightweight model adjustments, enabling flexible customization through different architectural modifications. Content-preserving methods [48, 63, 72, 24] emphasize maintaining subject identity and semantic consistency through specialized training strategies, introducing techniques for robust feature preservation and identity-aware generation. Advanced control mechanisms have emerged through attention-based techniques [15, 20, 1, 24, 14] and plug-and-play solutions [58], alongside lightweight training-free alternatives including editing-based approaches [3, 5, 61, 67, 41, 19, 7, 37, 9, 26] that enable flexible attribute manipulation through direct image editing and feature-level modifications without model retraining. Despite these developments, existing approaches face inherent limitations in achieving effective feature disentanglement while maintaining parameter efficiency, particularly when handling multiple customization objectives. These challenges stem from the complex nature of visual feature separation and the computational overhead required for precise attribute control, motivating our investigation into a more unified and parameter-efficient approach that leverages orthogonal basis vectors for disentanglement.
Feature disentanglement, particularly in the context of content and style separation, has been extensively studied in text-to-image models. Traditional approaches [48, 73, 38] focus on single-aspect control but struggle with joint manipulation of multiple attributes. Several methods have been proposed for merging multiple LoRA adaptations, such as Custom Diffusion [24], Mix-of-Show [12], Ortha [43], LoRACLR [51], they often require complex training strategies or additional regularization terms. In parallel, model-specific architectures [62, 49, 28] partially address this limitation through specialized designs for feature disentanglement, though their reliance on specific architectural choices limits their generality, with methods like B-LoRA [8] leveraging SDXL’s architectural characteristics to identify specific blocks crucial for style-content separation, and ZipLoRA [49] proposing efficient merging strategies for independently trained style and subject LoRAs. While these approaches demonstrate promising results within their targeted frameworks, they remain heavily dependent on model-specific architectures (particularly SDXL) and often face challenges in maintaining both subject and style fidelity during joint generation.
3 Method
3.1 Preliminaries
Continuous generative frameworks. Modern generative models often construct samples through learned continuous dynamics that bridge data distributions and tractable priors. Two prominent paradigms—diffusion models [16, 54, 47, 44] and flow matching [32, 30, 6, 25], share this philosophy but differ in their construction of probability paths and training objectives. Both frameworks leverage continuous transformations to map between a complex data distribution and a simple prior distribution, enabling efficient sampling and learning. Diffusion models gradually corrupt observed data $\mathbf{x}$ by mixing it with Gaussian noise ($\boldsymbol{\epsilon}$) over time: $\mathbf{z}_{t}=\alpha_{t}\mathbf{x}+\sigma_{t}\boldsymbol{\epsilon}$. The goal is to learn a reverse process that can reconstruct the original data from the noisy samples. Training involves estimating either $\hat{\mathbf{x}}$ or $\hat{\boldsymbol{\epsilon}}$ using a neural network:
$\mathcal{L}(\mathbf{x})=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[w(\lambda_{t})\cdot\frac{\mathrm{d}\lambda}{\mathrm{d}t}\cdot\|\boldsymbol{\hat{\epsilon}}-\boldsymbol{\epsilon}\|_{2}^{2}\right],$ (1)
where $\lambda_{t}=\log(\alpha_{t}^{2}/\sigma_{t}^{2})$ is the log signal-to-noise ratio, and $w(\lambda_{t})$ is a weighting function that balances the importance of different noise levels [22]. Flow matching views the forward process as a linear interpolation between the data $\mathbf{x}$ and a noise term $\boldsymbol{\epsilon}$: $\mathbf{z}_{t}=(1-t)\mathbf{x}+t\boldsymbol{\epsilon}$. The key idea is to learn a velocity field $\mathbf{u}_{t}$ that describes the continuous transformation from noise to data:
$\mathcal{L}_{\mathrm{CFM}}(\mathbf{x})=\mathbb{E}_{t\sim\mathcal{U}(0,1),\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\|\hat{\mathbf{u}}-\mathbf{u}\|_{2}^{2}\right],$ (2)
where $\hat{\mathbf{u}}$ is the predicted velocity field, which can be expressed as a linear combination of $\hat{\boldsymbol{\epsilon}}$ and $\mathbf{z}_{t}$. This objective can be rewritten in terms of the noise prediction, similar to Eq. 1, with an appropriate weighting function.
LoRA fine-tuning. LoRA has emerged as a parameter-efficient fine-tuning approach for adapting large pre-trained models to downstream tasks. Given a pre-trained weight matrix $W\in\mathbb{R}^{m\times n}$, LoRA learns a low-rank update $\Delta W=BA$, where $B\in\mathbb{R}^{m\times r}$ and $A\in\mathbb{R}^{r\times n}$ are trainable matrices with rank $r\ll\min(m,n)$. This decomposition significantly reduces the number of trainable parameters from $mn$ to $r(m+n)$. In text-to-image models [44, 6, 25], LoRA is typically injected into key components such as cross-attention and self-attention [59] layers, we provide detailed injection configurations in the Appendix. During training, the original objective functions (i.e., Eq. 1 or Eq. 2) remain unchanged, but the optimization is performed only on the low-rank parameters $A$ and $B$, while keeping the pretrained weights $W$ fixed, as shown in Figure 4 (b).
Figure 2: Maximum and minimum cosine similarities across different matrices. $Q$ and $R$ matrices are obtained from QR-LoRA framework with direct fine-tuning strategy, $\Delta R$ matrices from QR-LoRA framework with $\Delta R$-only update strategy, and $A$, $B$ matrices from vanilla LoRA decomposition. The analysis reveals distinct similarity patterns across these matrix types, demonstrating their unique roles in feature representation. All similarity values are computed between matrices trained on different image pairs. See Appendix for experimental cases and training details.
3.2 Motivation and Analysis
Existing methods like ZipLoRA [49], ComposLoRA [74], and CMLoRA [76], while maintaining the original LoRA [17] training paradigm, share a common approach: they first train independent LoRA modules with the standard structure and then focus on studying their merging strategies. However, is this paradigm the optimal approach? This raises a fundamental question: Could we redesign the LoRA structure from the ground up to achieve inherent feature disentanglement and better merging capabilities? Through extensive matrix similarity analysis, we identify two critical insights that motivate our approach:
Feature entanglement in direct weight updates. The conventional LoRA training paradigm and direct QR decomposition training both exhibit similar feature entanglement issues. Our analysis reveals high cosine similarities between task-specific updates (Figure 2), particularly in the matrices $A$ and $Q$. While the $B$ matrices show relatively lower similarities, the $R$ matrices demonstrate notably higher similarities ($>0.75$) due to our initialization strategy that inherits substantial prior information from the original weights (detailed in Algorithm 1 and Sec 3.3). This similarity pattern reveals a crucial insight: the stability of $Q$ matrices across different tasks suggests their potential as fixed orthogonal bases, while the $R$ matrices provides suitable flexibility for task-specific adaptations. Since both approaches require training multiple components (i.e., $A$,$B$ and $Q$,$R$), achieving clean feature separation during merging becomes challenging. This observation motivates our design choice to focus on learning task-specific updates $\Delta R$ rather than training the entire decomposition. The effectiveness of this approach is validated by our empirical results in Figure 2 and Figure 3, which show that the maximum cosine similarity between $\Delta R$ matrices across all injection layers remains consistently below 0.2, with mean values stable around 0, indicating superior feature disentanglement.
Figure 3: Comparison of cosine similarities between different training strategies in QR-LoRA. Visualization of layer-wise cosine similarities between matrices obtained from training on a randomly selected image pair in SDXL model. The comparison demonstrates distinct characteristics between directly fine-tuning $Q$ and $R$ matrices (upper zoom-in) versus only tuning $\Delta R$ matrices (lower zoom-in). See Appendix for comprehensive analyses across different model architectures and training data.
Lack of structural constraints. Most existing methods treat weight updates as unconstrained matrices, without leveraging mathematical properties that could facilitate feature disentanglement. Our investigation shows that imposing appropriate structural constraints through matrix decomposition can lead to more effective feature separation. Specifically, we find that orthogonal bases provide a natural framework for maintaining feature independence while allowing flexible combinations (see Appendix for detailed theoretical analysis). Inspired by these observations and the properties revealed in Figure 2 and Figure 3, we propose to leverage QR decomposition’s unique characteristics. The orthogonal matrix $Q$ provides an ideal basis for representing features independently, while the upper triangular matrix $R$ captures the essential transformations. This enables us to design a novel LoRA variant that naturally preserves feature disentanglement from the outset, rather than attempting to achieve it through post-training merging strategies.
Figure 4: Overview of QR-LoRA framework. Upper (Sec 3.3): Technical illustration of our method compared to traditional fine-tuning paradigms, highlighting our efficient parameter updates through orthogonal decomposition. Lower (Sec 3.4): Application pipeline of our framework for content-style disentanglement, consisting of a disentangled fine-tuning module and a streamlined feature merging module.
3.3 QR-LoRA Methodology
Based on our analysis of the limitations in existing approaches and the insights gained from matrix similarity studies, we propose QR-LoRA, a novel parameter-efficient fine-tuning framework that naturally facilitates feature disentanglement through orthogonal decomposition. The key innovation of our approach lies in its unique initialization strategy and training mechanism, which we detail below.
Core information extraction. As illustrated in the upper part of Figure 4, compared to traditional full fine-tuning and LoRA, our QR-LoRA achieves more efficient parameter updates through orthogonal decomposition. Given a pretrained weight matrix $W\in\mathbb{R}^{m\times n}$, we first employ SVD to systematically extract its core information structure (Figure 4 (c)). This decomposition provides a natural way to identify and isolate the most significant components of the weight matrix:
$W=U\Sigma V^{\top}=\sigma_{1}u_{1}v_{1}^{\top}+\sigma_{2}u_{2}v_{2}^{\top}+\dots+\sigma_{s}u_{s}v_{s}^{\top},$ (3)
where $s\!=\!min(m,n)$, $U\in\mathbb{R}^{m\times s}$ contains the left singular vectors, $\Sigma\in\mathbb{R}^{s\times s}$ is a diagonal matrix of singular values arranged in descending order ($\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{s}$), reflecting the decreasing importance of feature dimensions, and $V^{\top}\in\mathbb{R}^{s\times n}$ contains the right singular vectors. By leveraging this natural ordering where larger singular values correspond to more significant feature transformations, we can construct a rank-$r$ core matrix that captures the most essential feature transformations:
$W_{core}=U_{[:,:r]}\Sigma_{[:r]}V^{\top}_{[:r,:]}\triangleq\sigma_{1}u_{1}v_{1}^{\top}+\dots+\sigma_{r}u_{r}v_{r}^{\top},$ (4)
this strategic truncation yields two complementary components: the core matrix $W_{core}$ that encapsulates the dominant feature transformations, and a complement matrix $W_{comp}=W-W_{core}$ that retains the residual information. This decomposition serves a dual purpose: it ensures that our subsequent QR-based adaptation focuses on the most significant components while maintaining structural consistency with the original LoRA framework through the initialization of the pretrained weight matrix $W$ with $W_{comp}$, which is outlined in Algorithm 1 lines 1-4.
Orthogonal basis construction. To effectively capture and manipulate feature transformations, we leverage QR decomposition to construct orthogonal bases for our parameter updates. Rather than directly decomposing $W_{core}$, we strategically apply QR decomposition to its transpose $W_{core}^{\top}$, which aligns with our goal of maintaining consistent matrix multiplication order with the original LoRA formulation ($\Delta W=BA\Leftrightarrow R^{\top}Q^{\top}$). Given that $W_{core}^{\top}=V_{[:r,:]}\Sigma_{[:r]}U^{\top}_{[:,:r]}$, we can naturally express it as a product of two low-rank matrices ($S$ and $T$):
$W_{core}^{\top}$ $=\underbrace{V_{[:r,:]}\Sigma_{[:r]}}_{S}\>\ \underbrace{U^{\top}_{[:,:r]}}_{T}=(Q_{s}R_{s})T$ (5)
$=\underbrace{Q_{s}}_{Q}\underbrace{(R_{s}T)}_{R}\triangleq QR,$
where $S$ characterizes the core structure of the column space of $W_{core}^{\top}$. By performing reduced QR decomposition ($Q_{s}R_{s}$) on $S$ rather than the full matrix, we obtain more precise orthogonal bases while maintaining computational efficiency. Through initializing the task-specific update matrix $\Delta R$ to zero, the final initialization of our QR-LoRA framework takes the form:
$W_{origin}=W_{comp}+(Q(R+\Delta R))^{\top},$ (6)
where $W_{origin}$ represents the original pretrained weights. By keeping both $Q$ and $R$ fixed during training, we provide stable anchors for task-specific adaptations through $\Delta R$. The complete process is detailed in Algorithm 1.
3.4 Task-Specific Training
As illustrated in the lower part of Figure 4, we apply our QR-LoRA framework to the challenging task of content-style disentanglement in image generation. Our approach consists of two key components that leverage the inherent orthogonality properties of our framework:
Disentangled fine-tuning. Through efficient fine-tuning on individual images, we separately learn content-specific ($\Delta R_{c}$) and style-specific ($\Delta R_{s}$) update matrices. The orthogonal basis $Q$ provided by our framework naturally facilitates this disentanglement by ensuring that different feature transformations remain independent (i.e., $\Delta R_{c}\perp\Delta R_{s}$ as shown in Figure 2). This independence is further reinforced by our training strategy where each $\Delta R$ matrix is optimized for its specific task (e.g., content or style) while maintaining the shared orthogonal basis $Q$ and base matrix $R$ fixed.
Streamlined feature merging. The inherent disentanglement properties of our learned $\Delta R$ matrices enable a remarkably simple yet effective merging strategy. Unlike existing approaches that require complex merging mechanisms, our method achieves flexible content-style integration through straightforward element-wise addition of the update matrices:
$\Delta R_{cs}=\lambda_{c}\Delta R_{c}+\lambda_{s}\Delta R_{s}$ (7)
where $\lambda_{c}$ and $\lambda_{s}$ are scaling coefficients that provide fine-grained control over the contribution of content and style features in the final generation. By default, $\lambda_{c}=\lambda_{s}=1$.
1
Input: Pretrained weight matrix $W\in\mathbb{R}^{m\times n}$;
Target task type $\tau$ (content or style);
Rank $r$ for low-rank approximation; Learning rate $\eta$
Output: Trained $\Delta R_{\tau}$; Orthogonal basis $Q$;
Upper triangular base matrix $R$
2
$U,\Sigma,V^{\top}\leftarrow\textbf{SVD}(W)$
  // SVD of $W$
$W_{core}\leftarrow U_{[:,:r]}\Sigma_{[:r]}V^{\top}_{[:r,:]}$
  // Core matrix
3 $W_{comp}\leftarrow W-W_{core};W_{origin}\leftarrow W$
$W\leftarrow W_{comp}$
  // Complement matrix
4 $S\leftarrow V_{[:r,:]}\Sigma_{[:r]}\in\mathbb{R}^{n\times r};T\leftarrow U^{\top}_{[:,:r]}\in\mathbb{R}^{r\times m}$
$Q_{s},R_{s}\leftarrow\textbf{QR}(S);Q_{s}\in\mathbb{R}^{n\times r},R_{s}\in\mathbb{R}^{r\times r}$
  // Reduced QR decomposition
5
[…]
Initialize $\Delta R_{\tau}\leftarrow 0_{r\times m}$
  // Init to zero
7
[…]
9while not converged do
       $\mathcal{L}\leftarrow\textbf{TaskLoss}\{W_{comp}+(Q(R+\Delta R_{\tau}))^{\top}\}$ $\Delta R_{\tau}\leftarrow\Delta R_{\tau}-\eta\nabla_{\Delta R_{\tau}}\mathcal{L}$
        // Update
10
11
return $\Delta R_{\tau},Q,R$
Algorithm 1 QR-LoRA for Efficient and Disentangled Fine-tuning with Orthogonal Basis
4 Experiments
Figure 5: Qualitative comparison. Comparison of our QR-LoRA against state-of-the-art methods on SDXL and a naive baseline on SD3 and FLUX.1-dev models, demonstrating the model-agnostic nature and superior performance of our framework. Zoom in to view details.
Table 1: Quantitative comparison across different backbone models (SDXL, SD3, FLUX.1-dev). We report cosine similarities between generated images and their style (-S) or content (-C) references using DINO and CLIP features, along with user ratings (0-5 scale).
SDXL [44] SD3 [6] FLUX.1-dev [25]
ZipLoRA [49] B-LoRA [28] StyleAligned [15] Ours DB-LoRA [60] Ours DB-LoRA [60] Ours
DINO-S [2] $\uparrow$ 0.686 $\pm$ 0.131 0.689 $\pm$ 0.075 0.651 $\pm$ 0.061 0.694 $\pm$ 0.084 0.675 $\pm$ 0.043 0.711 $\pm$ 0.061 0.723 $\pm$ 0.105 0.744 $\pm$ 0.077
DINO-C [2] $\uparrow$ 0.712 $\pm$ 0.112 0.740 $\pm$ 0.064 0.758 $\pm$ 0.053 0.776 $\pm$ 0.065 0.673 $\pm$ 0.032 0.690 $\pm$ 0.040 0.654 $\pm$ 0.072 0.687 $\pm$ 0.086
CLIP-S [45] $\uparrow$ 0.686 $\pm$ 0.105 0.669 $\pm$ 0.085 0.658 $\pm$ 0.076 0.707 $\pm$ 0.060 0.652 $\pm$ 0.071 0.703 $\pm$ 0.056 0.641 $\pm$ 0.074 0.690 $\pm$ 0.069
CLIP-C [45] $\uparrow$ 0.668 $\pm$ 0.149 0.646 $\pm$ 0.098 0.662 $\pm$ 0.114 0.709 $\pm$ 0.087 0.771 $\pm$ 0.081 0.792 $\pm$ 0.077 0.677 $\pm$ 0.125 0.709 $\pm$ 0.096
User-Study $\uparrow$ 3.13 3.34 3.67 4.07 2.93 4.14 2.86 3.96
4.1 Implementation details
Experimental setup. To validate the model-agnostic nature of our framework, we evaluate QR-LoRA on multiple state-of-the-art diffusion models including SDXL [44], SD3 [6], and FLUX.1-dev [25]. During the fine-tuning process, we keep the pretrained model weights and text encoders frozen. All experiments are conducted on a single image per training instance. We employ the Adam [23] optimizer with a learning rate of $-41e\text{-}$ 4 without data augmentations. The LoRA rank is set to $r=64$ with 500 training steps, requiring approximately 20 minutes on a single A800 GPU. Our training data is sourced from previous works [48, 52, 65] and publicly available content. Following previous works [48, 52, 49], we adopt a standard prompt template for training. Specifically, we use $<\!\!c\!\!>$ to denote content-related trigger words (e.g., “[c]” or “sks”) and $<\!\!s\!\!>$ for style-related expressions (e.g., “[s]” or simple artistic description [49, 28]), combining them in the format “a $<\!\!c\!\!>$ in $<\!\!s\!\!>$ style”. For fair comparison, we maintain consistent content and style image settings across all comparative experiments.
Compared methods. For SDXL [44], where most recent style transfer methods are developed, we compare against state-of-the-art approaches including ZipLoRA [49], StyleAligned [15], and B-LoRA [28]. For SD3 [6] and FLUX.1-dev [25] models, we compare against DreamBooth-LoRA [60] using naive weight combination.
Metrics. To evaluate our method quantitatively, we compute cosine similarities between generated images and their references using DINO [2] and CLIP [45] features for both style (-S) and content (-C) aspects. We also conduct a user study to assess the perceptual quality, with detailed settings provided in the Appendix.
Figure 6: Visualization on different backbone models. Demonstration of QR-LoRA’s generation results across different backbone models, showing consistent high-quality performance.
4.2 Results and Analysis
We conduct comprehensive quantitative and qualitative experiments to evaluate our method, with additional experimental results provided in the supplementary material.
Qualitative Analysis. Figure 5 demonstrates the superior performance of QR-LoRA across various test scenarios. Our method consistently outperforms existing approaches in style transfer tasks, with particularly notable advantages in preserving human poses and anatomical details when handling complex body structures and dynamic postures. Moreover, QR-LoRA exhibits strong model-agnostic capabilities, achieving consistent high-quality results across different backbone models including SDXL, SD3, and FLUX.1-dev, which is also demonstrated in Figure 6. This cross-model effectiveness demonstrates that our approach to feature manipulation and style transfer can be successfully applied to diverse model architectures while maintaining structural integrity of the generated images.
Quantitative Analysis We evaluate our method on a test set of 64 randomly sampled generated images. As shown in Table 1, QR-LoRA outperforms state-of-the-art methods on SDXL in both content preservation and style transfer, with significantly higher user ratings in our user study. On SD3 and FLUX.1-dev, our method consistently improves over the naive baseline DB-LoRA [60] (i.e., direct merge), demonstrating its model-agnostic effectiveness.
Ablation. To further validate the robustness of our method, we experiment with different combinations of scaling coefficients $\lambda_{c}$ and $\lambda_{s}$ during feature composition. As shown in Figure 7, the consistent high-quality results across various coefficient pairs demonstrate that our approach is inherently robust, indicating that this scaling flexibility is a natural property rather than a requirement of our method.
Figure 7: Ablation study on scaling coefficient $\lambda_{c}$ and $\lambda_{s}$.
4.3 Application
Adaptive Training Strategies. Our framework provides flexible training configurations for different scenarios as shown in Figure 8. For feature disentanglement tasks, training only the $\Delta R$ matrices achieves comparable convergence to standard LoRA with better generalization. When faster convergence is needed, training both $Q$ and $R$ components simultaneously accelerates the process by leveraging pretrained structural information.
Figure 8: Convergence analysis. Comparison of training convergence between QR-LoRA and LoRA.
Versatile Feature Composition. Beyond style-content synthesis, our method demonstrates remarkable capability in disentangled composition of various image attributes. As shown in Figure 9, QR-LoRA effectively handles diverse feature combinations. This versatility highlights the method’s potential for broader applications.
Figure 9: Applications on multi-feature composition. Our method enables flexible combination of various image attributes, demonstrating its capability beyond style-content synthesis.
5 Discussion
In this work, we have proposed QR-LoRA, a novel parameter-efficient fine-tuning framework that achieves effective feature disentanglement through orthogonal decomposition. By leveraging SVD-QR based initialization and a $\Delta R$ mechanism in a fixed orthogonal space, our approach reduces trainable parameters while enabling precise feature control. The framework demonstrates robust performance across different architectures, offering a complementary solution for feature disentanglement tasks.
Limitation. While our method demonstrates strong performance, the incorporation of matrix decomposition operations during initialization introduces additional computational overhead. Specifically, this initialization process requires $\sim$1, 2, and 5 minutes for SDXL, SD3, and FLUX.1-dev respectively. However, this one-time cost during initialization does not affect subsequent inference efficiency.
Future Work. The model-agnostic nature of QR-LoRA enables its extension to other modalities such as 3D [46, 66] and video [71, 56, 69, 70, 75] customization. Furthermore, our framework naturally aligns with Mixture-of-Experts (MoE) [33] architectures when extended to multiple $\Delta R$ matrices, where studying the collaborative dynamics could enable more flexible and powerful control capabilities.
Appendix of QR-LoRA: Efficient and Disentangled Fine-tuning via QR Decomposition for Customized Generation
(Supplementary Material)
Appendix A Theoretical Analysis of QR-LoRA Parameter Disentanglement
A.1 Minimal Frobenius Norm Property
Lemma 1 (Minimal Frobenius Norm Property of QR-LoRA Parameterization).
Consider a weight matrix $W\in\mathbb{R}^{m\times n}$ with a low-rank update $\Delta W=Q\Delta R$, where $Q\in\mathbb{R}^{m\times r}$ is a column-orthogonal matrix (i.e., $Q^{\top}Q=I_{r}$), and $\Delta R\in\mathbb{R}^{r\times n}$ represents trainable parameters. For any rank-$r$ update $\Delta W$, the parameterization $\Delta W=Q\Delta R$ achieves the minimal Frobenius norm if and only if $Q$ is column-orthogonal, yielding:
$\|\Delta W\|_{F}=\|Q\Delta R\|_{F}=\|\Delta R\|_{F},$ (8)
with $\Delta R$ having a unique solution under orthogonal constraints.
Proof.
Leveraging the norm-preserving property of orthogonal matrices, for any column-orthogonal $Q$, we have:
$\|Q\Delta R\|_{F}$ $=\sqrt{\text{tr}(\Delta R^{\top}Q^{\top}Q\Delta R)}$ (9)
$=\sqrt{\text{tr}(\Delta R^{\top}\Delta R)}$
$=\|\Delta R\|_{F},$
consequently, the Frobenius norm of the parameterized update $\Delta W=Q\Delta R$ is equivalent to the norm of $\Delta R$. ∎
This equivalence has profound implications for our framework:
  • Minimal Intervention Principle: The uniqueness of the minimal norm solution reflects the optimal task-specific transformation. For any target weight update $\Delta W$, we seek its representation in the form of $Q\Delta R$, where $Q$ is column-orthogonal. Among multiple solutions satisfying the task objective, the minimal norm solution maximally preserves the model’s original generalization capabilities while introducing only necessary task-specific transformations.
  • Task-Specific Feature Encoding: The minimal norm property of $\Delta R$ prevents overfitting by incorporating only the minimal changes required for the specific task. This principle ensures that $\Delta R$ exclusively encodes information relevant to the current task, avoiding the incorporation of general features shared across different tasks.
A.2 Gradient Update Decoupling Analysis
Theorem 1 (Statistical Independence of Orthogonal Projections).
In a Hilbert space with orthogonal basis, projections along different feature directions are statistically independent. Specifically, for any two orthogonal directions $v_{i}$ and $v_{j}$ ($v_{i}^{\top}v_{j}=0$), the projections of a random vector $x$ onto these directions, denoted as $P_{i}(x)$ and $P_{j}(x)$ respectively, are statistically uncorrelated:
$E[P_{i}(x)P_{j}(x)]=0,$ (10)
where $E[\cdot]$ denotes the expectation operator.
In the context of Hilbert spaces with orthogonal bases, projections along different feature directions exhibit statistical independence. This fundamental property plays a crucial role in our QR-LoRA framework, where the $Q$ matrix defines a set of orthogonal basis vectors, and the $\Delta R$ matrix represents coordinates in this orthogonal basis system. The gradient update mechanism in our framework demonstrates how this orthogonality naturally leads to feature disentanglement. For a loss function $\mathcal{L}(W)$ with the parameterization $W=W_{\text{comp}}+Q(R+\Delta R)$, the gradient update mechanism exhibits the following key property:
The gradient of the loss with respect to $\Delta R$ is given by:
$\frac{\partial\mathcal{L}}{\partial\Delta R}=\frac{\partial\mathcal{L}}{\partial W}\cdot\frac{\partial W}{\partial\Delta R}=Q^{\top}\frac{\partial\mathcal{L}}{\partial W}.$ (11)
Given the column orthogonality of $Q$ ($Q^{\top}Q=I$), this gradient formulation reveals that $\partial\mathcal{L}/\partial\Delta R$ is an orthogonal projection of $\partial\mathcal{L}/\partial W$ onto the column space of $Q$. This projection mechanism is particularly significant when training for different tasks (e.g., style or content adaptation), as these tasks inherently optimize distinct visual attributes. The orthogonality of their gradient directions in the high-dimensional space is preserved and enhanced through the $Q$ matrix projection, causing different $\Delta R$ matrices to evolve along approximately orthogonal trajectories during training.
To provide a comprehensive understanding of QR-LoRA’s theoretical advantages, we further analyze its forward pass and gradient computation mechanisms in comparison with existing methods. Table 2 presents a detailed comparison of the forward pass and gradient computation between QR-LoRA and other representative approaches.
Table 2: Summary of forward pass and gradient computation in different methods. This comparison highlights QR-LoRA’s unique orthogonal projection-based gradient update mechanism, which naturally leads to more effective feature disentanglement compared to traditional approaches.
LoRA [17] Ortha [43] QR-LoRA
Forward $Y\!\!=\!\!X(W\!\!+\!\!BA)$ $Y\!\!=\!\!X(W\!\!+\!\!BA)$ $Y=XW;W\!=\!U\Sigma V^{\top}$
$A\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{r\times n}$ $A\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{r\times n}$ $W_{core}^{\top}\!=\!V_{[:r,:]}\Sigma_{[:r]}U_{[:,:r]}^{\top}\!=\!QR$
$B\!=\!0\!\in\!\mathbb{R}^{m\times r}$ $B\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{m\times r}$ $W\!=\!W_{comp}+(Q(R+\Delta R))^{\top}$
Gradient $\frac{\partial\mathcal{L}}{\partial A}\!\!=\!\!B^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{ 0}$ frozen $B$, train $A$ $Q\!\!\in\!\!\mathbb{R}^{n\times r}\!,\!R\!\!\in\!\!\mathbb{R}^{r\times m}$; train $\Delta R$
$\frac{\partial\mathcal{L}}{\partial B}\!\!=\!\!\frac{\partial\mathcal{L}}{\partial W}\!A^{\top}\!\!\Rightarrow\!\textbf{ Random}$ $\frac{\partial\mathcal{L}}{\partial A}\!\!=\!\!B^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{Random}$ $\frac{\partial\mathcal{L}}{\partial\Delta R}\!\!=\!\!Q^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{Primary}$
As shown in Table 2, QR-LoRA fundamentally differs from previous approaches in both its forward pass computation and gradient update mechanism. While traditional LoRA [17] and Ortha [43] rely on random initialization and direct parameter updates, our method leverages orthogonal decomposition to achieve more principled and effective parameter updates.
A.3 Theoretical Interpretation of Experimental Observations
  • Stability of $Q$: The stability of $Q$ matrices across different tasks stems from their construction through SVD-based QR decomposition, where their directions are primarily determined by the underlying data distribution, leading to consistently high similarities.
  • Independence of $\Delta R$: The orthogonal projection property ensures that $\Delta R$ matrices exclusively capture task-specific information, resulting in significant variations that reflect the unique characteristics of different tasks.
Figure 10: Rank sensitivity analysis. Our method maintains stable performance across different rank settings, demonstrating its robustness to this hyperparameter choice.
Appendix B Additional Implementation Details
B.1 Implementation Details of Compared Methods.
In our comprehensive evaluation, we implement ZipLoRA based on an unofficial implementation [^1] https://github.com/mkshing/ziplora-pytorch, configuring it with rank $r=64$ and learning rate $-55e\text{-}$ 5. For StyleAligned and B-LoRA, we adopt their respective default configurations from the original papers. Additionally, we employ the standard DreamBooth-LoRA implementation for experiments on SD3 and FLUX.1-dev models, maintaining consistent hyperparameters ($r=64$, lr=$1e\text{-}4$) across all comparisons to ensure fairness.
Model-specific Training Configurations. Across all backbone models, we exclusively inject QR-LoRA modules into the main generative networks (e.g., UNet, MM-DiT) while keeping the text encoders frozen during training. For different backbone models, we inject QR-LoRA modules into specific layers as shown in Table 3:
  • SDXL: We inject our QR-LoRA modules into the self-attention layers of UNet only, focusing on the core architecture components. For inference, we use 50 denoising steps while keeping other parameters at their default values.
  • SD3: Following the model architecture, we apply our method to both standard attention blocks and additional projection layers unique to SD3. During inference, we employ 28 denoising steps with default configurations for other parameters.
  • FLUX.1-dev: We extend the injection to include both attention mechanisms and feedforward networks, covering the expanded architecture of FLUX.1-dev. Similar to SD3, we use 28 denoising steps for inference while maintaining default settings for other parameters.
Table 3: Injection layer types of QR-LoRA across different backbone models. This table details the specific architectural components where our QR-LoRA modules are injected during training across different foundation models (SDXL, SD3, and FLUX.1-dev), demonstrating the practical implementation of our method’s disentangled fine-tuning approach.
SDXL SD3 FLUX.1-dev
to_k attn.to_k attn.to_k
to_q attn.to_q attn.to_q
to_v attn.to_v attn.to_v
to_out.0 attn.to_out.0 attn.to_out.0
attn.add_k_proj attn.add_k_proj
attn.add_q_proj attn.add_q_proj
attn.add_v_proj attn.add_v_proj
attn.to_add_out attn.to_add_out
ff.net.0.proj
ff.net.2
ff_context.net.0.proj
ff_context.net.2
Appendix C Additional Analysis Details and Results
C.1 Experimental Setup for Matrix Similarity Analysis
To comprehensively evaluate the feature disentanglement capabilities of our QR-LoRA framework, we conduct extensive experiments using 27 randomly selected image pairs (sample index from 0-26), as visualized in Figure 13. For each image pair, we analyze three different training strategies: (1) a QR decomposition variant where both $Q$ and $R$ matrices are directly fine-tuned, (2) a $\Delta R$-only update strategy, and (3) vanilla LoRA with $A$ and $B$ matrices. The cosine similarities between corresponding matrices across different training instances are computed and recorded in Table 4, with their distributions visualized in Figure 2 of the main text. This experimental design allows us to analyze the inherent properties of different matrices and their roles in feature representation.
C.2 Cross-Model Analysis of Matrix Properties
We extend our analysis across three state-of-the-art text-to-image models: SDXL, SD3, and FLUX.1-dev, to validate the model-agnostic nature of our approach. As shown in Figure 14, Figure 16, and Figure 15, the layer-wise cosine similarity patterns remain consistent across different model architectures, demonstrating the generalizability of our QR-LoRA framework. Specifically:
  • $Q$ Matrix Properties: Across all models, $Q$ matrices maintain remarkably high similarities (0.90-0.99), indicating their role as stable orthogonal bases for feature representation.
  • $R$ Matrix Characteristics: The directly fine-tuned $R$ matrices show moderate but significant similarities (0.79-0.89), suggesting their inherited structural information from the original weights.
  • $\Delta R$ Matrix Behavior: In contrast, $\Delta R$ matrices consistently exhibit substantially lower similarities (-0.17-0.25), validating their effectiveness in capturing task-specific features independently.
Figure 11: User study interface. Screenshot of our evaluation interface where participants rate generated images on a 0-5 scale based on content preservation and style transfer quality.
C.3 Statistical Analysis of Matrix Similarities
The numerical results presented in Table 4 provide a detailed quantitative perspective on the similarity patterns. The data reveals several key insights:
  • The maximum similarity values for $Q$ matrices consistently approach 1.0 (0.9998-0.9999), with minimum values remaining above 0.90, demonstrating the stability of the orthogonal basis.
  • $R$ matrices show maximum similarities above 0.99 but with lower minimum values (0.79-0.89), reflecting their role in capturing task-relevant transformations while maintaining some degree of shared structure.
  • $\Delta R$ matrices exhibit significantly lower similarity values, with maximums below 0.26 and minimums reaching -0.17, quantitatively confirming their capacity for independent feature representation.
  • In comparison, traditional LoRA’s $A$ matrices show consistently high similarities (0.77-1.0), while $B$ matrices display moderate to low similarities (0.15-0.38), highlighting the advantages of our orthogonal decomposition approach.
This comprehensive analysis validates the theoretical foundations of our QR-LoRA framework and demonstrates its effectiveness in achieving feature disentanglement across different model architectures. The consistent patterns observed in both layer-wise visualizations and statistical distributions support our design choice of focusing on $\Delta R$ updates for efficient and clean feature separation.
Figure 12: Additional visualization results on different backbone models. Extended demonstration of QR-LoRA’s generation results across SDXL, SD3, and FLUX.1-dev models, further validating the consistent high-quality performance and model-agnostic nature of our approach. These supplementary examples provide a more comprehensive view of our method’s capabilities across diverse model architectures, complementing the results presented in Figure 6.
Figure 13: Detailed experimental setup for matrix similarity comparison in Figure 2. We conduct experiments on 27 randomly selected image pairs to analyze the properties of different matrices in our QR-LoRA framework. For each image in the pairs, we train two types of LoRA adaptations: one using QR decomposition where both $Q$ and $R$ matrices are directly fine-tuned and another using the $\Delta R$ formulation. We then compute the cosine similarities between $Q$ matrices, $R$ matrices, and $\Delta R$ matrices across all injection layers in the SDXL model.
Table 4: Numerical results of matrix similarity comparison in Figure 2. This table presents the detailed cosine similarity values for $Q$ matrices, $R$ matrices, $\Delta R$ matrices, $A$ matrices, and $B$ matrices from 27 pairs of image training data. For each image pair (i.e., sample index), we compute the cosine similarities across all injection layers in the SDXL model and record both the maximum ($Q_{max}$, $R_{max}$, $\Delta R_{max}$, $A_{max}$, $B_{max}$) and minimum ($Q_{min}$, $R_{min}$, $\Delta R_{min}$, $A_{min}$, $B_{min}$) similarity values.
Sample Index $Q_{max}$ $Q_{min}$ $R_{max}$ $R_{min}$ $\Delta R_{max}$ $\Delta R_{min}$ $A_{max}$ $A_{min}$ $B_{max}$ $B_{min}$
0 0.9999 0.9787 0.9997 0.8409 0.1657 -0.1139 1.0000 0.8046 0.3615 -0.0355
1 0.9999 0.9786 0.9997 0.8253 0.1632 -0.1223 1.0000 0.7970 0.3249 -0.0625
2 0.9999 0.9682 0.9997 0.8391 0.1135 -0.1038 1.0000 0.8161 0.3198 -0.0826
3 0.9998 0.9311 0.9996 0.8104 0.1959 -0.1335 1.0000 0.8075 0.2322 -0.0565
4 0.9998 0.9313 0.9996 0.7993 0.1241 -0.0818 1.0000 0.8020 0.2295 -0.0626
5 0.9998 0.9055 0.9997 0.8143 0.1529 -0.0872 1.0000 0.8144 0.2226 -0.0558
6 0.9998 0.9679 0.9997 0.8235 0.1566 -0.0581 1.0000 0.7917 0.2113 -0.0807
7 0.9999 0.9680 0.9997 0.8129 0.1421 -0.0699 1.0000 0.8126 0.2145 -0.0783
8 0.9999 0.9761 0.9997 0.8187 0.1257 -0.0913 1.0000 0.8053 0.1898 -0.2033
9 0.9999 0.9709 0.9998 0.8441 0.1749 -0.1036 1.0000 0.8272 0.3044 -0.1188
10 0.9999 0.9741 0.9998 0.8245 0.1500 -0.0783 1.0000 0.8163 0.2070 -0.0793
11 0.9999 0.9792 0.9998 0.8833 0.1399 -0.0611 1.0000 0.8379 0.1525 -0.1278
12 0.9999 0.9765 0.9997 0.8485 0.1816 -0.1094 1.0000 0.8253 0.2499 -0.0961
13 0.9999 0.9793 0.9998 0.8216 0.2493 -0.0463 1.0000 0.8164 0.1490 -0.1374
14 0.9999 0.9854 0.9999 0.8894 0.2067 -0.0933 1.0000 0.8332 0.1817 -0.1316
15 0.9998 0.9724 0.9996 0.8410 0.1756 -0.1686 1.0000 0.8416 0.3820 -0.0801
16 0.9998 0.9752 0.9996 0.8186 0.1851 -0.1050 1.0000 0.8361 0.2329 -0.0636
17 0.9998 0.9807 0.9997 0.8818 0.1262 -0.0695 1.0000 0.8413 0.2125 -0.1462
18 0.9999 0.9055 0.9996 0.8340 0.2539 -0.0760 1.0000 0.8346 0.3847 -0.2045
19 0.9999 0.9054 0.9997 0.8408 0.1209 -0.0420 1.0000 0.8398 0.2905 -0.0762
20 0.9999 0.9053 0.9996 0.8339 0.1705 -0.0540 1.0000 0.8298 0.2734 -0.0548
21 0.9998 0.9045 0.9995 0.8040 0.2042 -0.0466 1.0000 0.7727 0.1746 -0.0739
22 0.9998 0.9044 0.9995 0.8095 0.1441 -0.0716 1.0000 0.7714 0.1531 -0.1031
23 0.9998 0.9045 0.9995 0.7992 0.1473 -0.0588 1.0000 0.7853 0.1558 -0.0861
24 0.9998 0.9055 0.9996 0.8272 0.1711 -0.0512 1.0000 0.8034 0.1969 -0.0806
25 0.9998 0.9053 0.9996 0.8207 0.1686 -0.0851 1.0000 0.8147 0.2685 -0.0401
26 0.9998 0.9053 0.9996 0.8134 0.1719 -0.0850 1.0000 0.8108 0.2691 -0.0488
Figure 14: Extended analysis of cosine similarities under SDXL model, complementing Figure 3. We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the SDXL backbone model to validate the generalization of our findings. The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results. Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns. This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
Figure 15: Extended analysis of cosine similarities under FLUX.1-dev model, complementing Figure 3. We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the FLUX.1-dev backbone model to validate the generalization of our findings. The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results. Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns. This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
Figure 16: Extended analysis of cosine similarities under SD3 model, complementing Figure 3. We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the SD3 backbone model to validate the generalization of our findings. The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results. Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns. This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
Appendix D Additional Experimental Results and Discussion
D.1 Ablation Studies
Scaling coefficient. To comprehensively evaluate the robustness and flexibility of our QR-LoRA framework, we conduct extensive ablation studies on scaling coefficients across three different backbone models: SDXL, SD3, and FLUX.1-dev. Our experiments systematically explore coefficient combinations (i.e., $\lambda_{c},\lambda_{s}$) ranging from 0.5 to 1.0 with intervals of 0.1, resulting in 36 distinct combinations for each model. This granular analysis reveals several key insights about our method’s behavior and capabilities.
As demonstrated in Figures 17, 18, and 19, our approach maintains consistent generation quality across different coefficient settings while enabling fine-grained control over the content-style fusion process. The smooth transitions observed between different coefficient combinations indicate that our QR decomposition effectively disentangles content and style features, allowing for intuitive and stable interpolation between these attributes. Notably, this behavior remains consistent across all three model architectures, validating the model-agnostic nature of our approach.
Furthermore, our experiments demonstrate that our method exhibits robust performance across a wide range of scaling coefficients after obtaining the disentangled deltaR matrices. This scaling exploration serves as a deeper investigation into our method’s capabilities, rather than a limitation. It’s worth noting that such scaling strategies are generally applicable to any LoRA-based methods, showcasing the versatility of our approach in achieving various levels of content preservation and style transfer effects.
Rank. To comprehensively evaluate the robustness of our method, we conduct a sensitivity analysis across different rank settings. As shown in Figure 10, our method demonstrates consistent performance and stable generation quality across various rank values, indicating its inherent robustness to this hyperparameter. We adopt rank=64 for all experiments in the main paper, as it strikes an optimal balance between computational efficiency and generation quality. This choice aligns well with common practices in the literature [49, 8] and proves to be a reliable setting for most applications.
Figure 17: Comprehensive ablation study on scaling coefficients for content-style fusion based on SDXL. We demonstrate the generation results under various combinations of content and style scaling coefficients. The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations. Please zoom in to view details.
Figure 18: Comprehensive ablation study on scaling coefficients for content-style fusion based on SD3. We demonstrate the generation results under various combinations of content and style scaling coefficients. The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations. Please zoom in to view details.
Figure 19: Comprehensive ablation study on scaling coefficients for content-style fusion based on FLUX.1-dev. The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations. Please zoom in to view details.
D.2 User Study
To evaluate the perceptual quality of our results, we conducted a comprehensive user study with 47 participants (40 males and 7 females). The participants represented a diverse demographic, with ages ranging from 22 to 45 years (mean age: 31.2). Among the participants, 15 were computer vision researchers, 12 were professional designers, and 20 were general users with basic knowledge of image editing. This diverse composition ensured a balanced evaluation from both technical and aesthetic perspectives.
Each participant was asked to rate the quality of generated images from different methods on a scale of 0-5, where 0 indicates poor quality and 5 indicates excellent quality. To facilitate the evaluation process, we designed an intuitive rating interface (shown in Figure 11) that allows participants to easily assess and score the generated images. The evaluation criteria focused on both content preservation and style transfer aspects. The ratings reported in Table 1 represent the average scores given by participants for each method. The consistently higher ratings achieved by our method (4.07 for SDXL, 4.14 for SD3, and 3.96 for FLUX.1-dev) compared to baseline approaches demonstrate the superior perceptual quality across different backbone models.
D.3 Additional Qualitative Results
To comprehensively demonstrate the effectiveness and versatility of our QR-LoRA framework, we present extensive qualitative results across different backbone models and diverse application scenarios. First, we conduct a thorough comparative analysis through Figure 20, which showcases our method’s superior performance against state-of-the-art approaches across SDXL, SD3, and FLUX.1-dev models. The results consistently demonstrate QR-LoRA’s ability to achieve better content preservation while maintaining high-fidelity style transfer, regardless of the underlying model architecture.
Figure 20: Extended qualitative comparison results. Comparison of QR-LoRA against state-of-the-art methods on SDXL, SD3, and FLUX.1-dev models, demonstrating consistent superior performance across different scenarios. Please zoom in to view details.
To further validate the model-agnostic nature of our approach, we present a comprehensive set of generation results across different backbone architectures. Figures 21, 22, and 23 showcase diverse generation results on SDXL, demonstrating our method’s capability in handling various challenging scenarios while maintaining high visual quality and style fidelity. The results span a wide range of content types and artistic styles, from intricate architectural details to subtle texture variations.
Figure 21: Diverse generation results based on SDXL. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations. Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics. Please zoom in to view details.
Figure 22: Diverse generation results on SDXL. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations. Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics. Please zoom in to view details.
Figure 23: Diverse generation results on SDXL. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations. Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics. Please zoom in to view details.
Extending our analysis to other backbone architectures, Figure 24 presents generation results on SD3, while Figure 25 shows results using FLUX.1-dev as the backbone. These examples demonstrate that QR-LoRA maintains its effectiveness across different model architectures, consistently producing high-quality results that effectively balance content preservation and style transfer.
Figure 24: Diverse generation results on SD3. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
Figure 25: Diverse generation results on FLUX.1-dev. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations. Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics. Please zoom in to view details.
Collectively, these results validate the robustness and generalization capability of our approach across diverse scenarios. The consistent high-quality performance across different content types (e.g., portraits, objects) and artistic styles (e.g., watercolor, digital art) demonstrates QR-LoRA’s effectiveness in handling complex geometric structures, fine-grained textures, and dramatic style variations. This comprehensive evaluation further supports the model-agnostic nature of our method and its potential for broad application in various image generation and style transfer tasks.
D.4 Ability to Re-contextualize
Our method not only excels at fusing content and style but also enables further customization through text prompts. While maintaining the high-quality integration of content and style elements, our approach supports flexible scene manipulation through additional textual instructions, as shown in Figure 26. For instance, given a content image of a dog and a specific artistic style, we can generate variations like ”wearing a hat” or ”riding a bicycle” while preserving the desired style characteristics. This re-contextualization capability is particularly valuable in practical artistic applications, as it allows artists to explore diverse creative possibilities while maintaining consistent stylistic elements across different scenarios.
Figure 26: Demonstration of our method’s re-contextualization capability. Given a content image and a style reference, our approach can generate variations through additional text prompts while maintaining style consistency. The examples show how different textual instructions (e.g., ”wearing a hat”, ”riding a bicycle”) guide the generation process to create diverse scenarios while preserving the original style characteristics.
D.5 The End
We anticipate that our model-agnostic, low-rank efficient fine-tuning approach will provide a versatile solution for fusion tasks requiring robust disentanglement properties, such as the fusion of arbitrary style and content. This stands in contrast to the field’s ongoing trend of pursuing specific disentanglement properties within particular models.

### v2

@@
 Experiments demonstrate that QR-LoRA achieves superior disentanglement in content-style fusion tasks, establishing a new paradigm for parameter-efficient, disentangled fine-tuning in generative models.
+The project page is available at: https://luna-ai-lab.github.io/QR-LoRA/
 Figure 1: Given a single content and a single style image, we present QR-LoRA, a method to achieve efficient and disentangled control of content and style features through orthogonal decomposition.
@@
 Following previous works [48, 52, 49], we adopt a standard prompt template for training.
-Specifically, we use $<\!\!c\!\!>$ to denote content-related trigger words (e.g., “[c]” or “sks”) and $<\!\!s\!\!>$ for style-related expressions (e.g., “[s]” or simple artistic description [49, 28]), combining them in the format “a $<\!\!c\!\!>$ in $<\!\!s\!\!>$ style”.
+Specifically, we use $<\!\!c\!\!>$ to denote content-related trigger words (e.g., “[c]" or “sks") and $<\!\!s\!\!>$ for style-related expressions (e.g., “[s]" or simple artistic description [49, 28]), combining them in the format “a $<\!\!c\!\!>$ in $<\!\!s\!\!>$ style".
 For fair comparison, we maintain consistent content and style image settings across all comparative experiments.
@@
 5 Discussion
-In this work, we have proposed QR-LoRA, a novel parameter-efficient fine-tuning framework that achieves effective feature disentanglement through orthogonal decomposition.
-By leveraging SVD-QR based initialization and a $\Delta R$ mechanism in a fixed orthogonal space, our approach reduces trainable parameters while enabling precise feature control.
-The framework demonstrates robust performance across different architectures, offering a complementary solution for feature disentanglement tasks.
 Limitation.
@@
 Furthermore, our framework naturally aligns with Mixture-of-Experts (MoE) [33] architectures when extended to multiple $\Delta R$ matrices, where studying the collaborative dynamics could enable more flexible and powerful control capabilities.
-Appendix of QR-LoRA: Efficient and Disentangled Fine-tuning via QR Decomposition for Customized Generation
-(Supplementary Material)
-Appendix A Theoretical Analysis of QR-LoRA Parameter Disentanglement
-A.1 Minimal Frobenius Norm Property
-Lemma 1 (Minimal Frobenius Norm Property of QR-LoRA Parameterization).
-Consider a weight matrix $W\in\mathbb{R}^{m\times n}$ with a low-rank update $\Delta W=Q\Delta R$, where $Q\in\mathbb{R}^{m\times r}$ is a column-orthogonal matrix (i.e., $Q^{\top}Q=I_{r}$), and $\Delta R\in\mathbb{R}^{r\times n}$ represents trainable parameters.
-For any rank-$r$ update $\Delta W$, the parameterization $\Delta W=Q\Delta R$ achieves the minimal Frobenius norm if and only if $Q$ is column-orthogonal, yielding:
-$\|\Delta W\|_{F}=\|Q\Delta R\|_{F}=\|\Delta R\|_{F},$ (8)
-with $\Delta R$ having a unique solution under orthogonal constraints.
-Proof.
-Leveraging the norm-preserving property of orthogonal matrices, for any column-orthogonal $Q$, we have:
-$\|Q\Delta R\|_{F}$ $=\sqrt{\text{tr}(\Delta R^{\top}Q^{\top}Q\Delta R)}$ (9)
-$=\sqrt{\text{tr}(\Delta R^{\top}\Delta R)}$
-$=\|\Delta R\|_{F},$
-consequently, the Frobenius norm of the parameterized update $\Delta W=Q\Delta R$ is equivalent to the norm of $\Delta R$. ∎
-This equivalence has profound implications for our framework:
-  • Minimal Intervention Principle: The uniqueness of the minimal norm solution reflects the optimal task-specific transformation.
-For any target weight update $\Delta W$, we seek its representation in the form of $Q\Delta R$, where $Q$ is column-orthogonal.
-Among multiple solutions satisfying the task objective, the minimal norm solution maximally preserves the model’s original generalization capabilities while introducing only necessary task-specific transformations.
-  • Task-Specific Feature Encoding: The minimal norm property of $\Delta R$ prevents overfitting by incorporating only the minimal changes required for the specific task.
-This principle ensures that $\Delta R$ exclusively encodes information relevant to the current task, avoiding the incorporation of general features shared across different tasks.
-A.2 Gradient Update Decoupling Analysis
-Theorem 1 (Statistical Independence of Orthogonal Projections).
-In a Hilbert space with orthogonal basis, projections along different feature directions are statistically independent.
-Specifically, for any two orthogonal directions $v_{i}$ and $v_{j}$ ($v_{i}^{\top}v_{j}=0$), the projections of a random vector $x$ onto these directions, denoted as $P_{i}(x)$ and $P_{j}(x)$ respectively, are statistically uncorrelated:
-$E[P_{i}(x)P_{j}(x)]=0,$ (10)
-where $E[\cdot]$ denotes the expectation operator.
-In the context of Hilbert spaces with orthogonal bases, projections along different feature directions exhibit statistical independence.
-This fundamental property plays a crucial role in our QR-LoRA framework, where the $Q$ matrix defines a set of orthogonal basis vectors, and the $\Delta R$ matrix represents coordinates in this orthogonal basis system.
-The gradient update mechanism in our framework demonstrates how this orthogonality naturally leads to feature disentanglement.
-For a loss function $\mathcal{L}(W)$ with the parameterization $W=W_{\text{comp}}+Q(R+\Delta R)$, the gradient update mechanism exhibits the following key property:
-The gradient of the loss with respect to $\Delta R$ is given by:
-$\frac{\partial\mathcal{L}}{\partial\Delta R}=\frac{\partial\mathcal{L}}{\partial W}\cdot\frac{\partial W}{\partial\Delta R}=Q^{\top}\frac{\partial\mathcal{L}}{\partial W}.$ (11)
-Given the column orthogonality of $Q$ ($Q^{\top}Q=I$), this gradient formulation reveals that $\partial\mathcal{L}/\partial\Delta R$ is an orthogonal projection of $\partial\mathcal{L}/\partial W$ onto the column space of $Q$.
-This projection mechanism is particularly significant when training for different tasks (e.g., style or content adaptation), as these tasks inherently optimize distinct visual attributes.
-The orthogonality of their gradient directions in the high-dimensional space is preserved and enhanced through the $Q$ matrix projection, causing different $\Delta R$ matrices to evolve along approximately orthogonal trajectories during training.
-To provide a comprehensive understanding of QR-LoRA’s theoretical advantages, we further analyze its forward pass and gradient computation mechanisms in comparison with existing methods.
-Table 2 presents a detailed comparison of the forward pass and gradient computation between QR-LoRA and other representative approaches.
-Table 2: Summary of forward pass and gradient computation in different methods.
-This comparison highlights QR-LoRA’s unique orthogonal projection-based gradient update mechanism, which naturally leads to more effective feature disentanglement compared to traditional approaches.
-LoRA [17] Ortha [43] QR-LoRA
-Forward $Y\!\!=\!\!X(W\!\!+\!\!BA)$ $Y\!\!=\!\!X(W\!\!+\!\!BA)$ $Y=XW;W\!=\!U\Sigma V^{\top}$
-$A\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{r\times n}$ $A\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{r\times n}$ $W_{core}^{\top}\!=\!V_{[:r,:]}\Sigma_{[:r]}U_{[:,:r]}^{\top}\!=\!QR$
-$B\!=\!0\!\in\!\mathbb{R}^{m\times r}$ $B\!\!\sim\!\!\mathcal{N}\!(\!0,\!\sigma^{2}\!)\!\!\in\!\!\mathbb{R}^{m\times r}$ $W\!=\!W_{comp}+(Q(R+\Delta R))^{\top}$
-Gradient $\frac{\partial\mathcal{L}}{\partial A}\!\!=\!\!B^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{ 0}$ frozen $B$, train $A$ $Q\!\!\in\!\!\mathbb{R}^{n\times r}\!,\!R\!\!\in\!\!\mathbb{R}^{r\times m}$; train $\Delta R$
-$\frac{\partial\mathcal{L}}{\partial B}\!\!=\!\!\frac{\partial\mathcal{L}}{\partial W}\!A^{\top}\!\!\Rightarrow\!\textbf{ Random}$ $\frac{\partial\mathcal{L}}{\partial A}\!\!=\!\!B^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{Random}$ $\frac{\partial\mathcal{L}}{\partial\Delta R}\!\!=\!\!Q^{\top}\!\!\frac{\partial\mathcal{L}}{\partial W}\!\!\Rightarrow\!\textbf{Primary}$
-As shown in Table 2, QR-LoRA fundamentally differs from previous approaches in both its forward pass computation and gradient update mechanism.
-While traditional LoRA [17] and Ortha [43] rely on random initialization and direct parameter updates, our method leverages orthogonal decomposition to achieve more principled and effective parameter updates.
-A.3 Theoretical Interpretation of Experimental Observations
-  • Stability of $Q$: The stability of $Q$ matrices across different tasks stems from their construction through SVD-based QR decomposition, where their directions are primarily determined by the underlying data distribution, leading to consistently high similarities.
-  • Independence of $\Delta R$: The orthogonal projection property ensures that $\Delta R$ matrices exclusively capture task-specific information, resulting in significant variations that reflect the unique characteristics of different tasks.
-Figure 10: Rank sensitivity analysis. Our method maintains stable performance across different rank settings, demonstrating its robustness to this hyperparameter choice.
-Appendix B Additional Implementation Details
-B.1 Implementation Details of Compared Methods.
-In our comprehensive evaluation, we implement ZipLoRA based on an unofficial implementation [^1] https://github.com/mkshing/ziplora-pytorch, configuring it with rank $r=64$ and learning rate $-55e\text{-}$ 5.
-For StyleAligned and B-LoRA, we adopt their respective default configurations from the original papers.
-Additionally, we employ the standard DreamBooth-LoRA implementation for experiments on SD3 and FLUX.1-dev models, maintaining consistent hyperparameters ($r=64$, lr=$1e\text{-}4$) across all comparisons to ensure fairness.
-Model-specific Training Configurations.
-Across all backbone models, we exclusively inject QR-LoRA modules into the main generative networks (e.g., UNet, MM-DiT) while keeping the text encoders frozen during training.
-For different backbone models, we inject QR-LoRA modules into specific layers as shown in Table 3:
-  • SDXL: We inject our QR-LoRA modules into the self-attention layers of UNet only, focusing on the core architecture components.
-For inference, we use 50 denoising steps while keeping other parameters at their default values.
-  • SD3: Following the model architecture, we apply our method to both standard attention blocks and additional projection layers unique to SD3.
-During inference, we employ 28 denoising steps with default configurations for other parameters.
-  • FLUX.1-dev: We extend the injection to include both attention mechanisms and feedforward networks, covering the expanded architecture of FLUX.1-dev.
-Similar to SD3, we use 28 denoising steps for inference while maintaining default settings for other parameters.
-Table 3: Injection layer types of QR-LoRA across different backbone models.
-This table details the specific architectural components where our QR-LoRA modules are injected during training across different foundation models (SDXL, SD3, and FLUX.1-dev), demonstrating the practical implementation of our method’s disentangled fine-tuning approach.
-SDXL SD3 FLUX.1-dev
-to_k attn.to_k attn.to_k
-to_q attn.to_q attn.to_q
-to_v attn.to_v attn.to_v
-to_out.0 attn.to_out.0 attn.to_out.0
-attn.add_k_proj attn.add_k_proj
-attn.add_q_proj attn.add_q_proj
-attn.add_v_proj attn.add_v_proj
-attn.to_add_out attn.to_add_out
-ff.net.0.proj
-ff.net.2
-ff_context.net.0.proj
-ff_context.net.2
-Appendix C Additional Analysis Details and Results
-C.1 Experimental Setup for Matrix Similarity Analysis
-To comprehensively evaluate the feature disentanglement capabilities of our QR-LoRA framework, we conduct extensive experiments using 27 randomly selected image pairs (sample index from 0-26), as visualized in Figure 13.
-For each image pair, we analyze three different training strategies: (1) a QR decomposition variant where both $Q$ and $R$ matrices are directly fine-tuned, (2) a $\Delta R$-only update strategy, and (3) vanilla LoRA with $A$ and $B$ matrices.
-The cosine similarities between corresponding matrices across different training instances are computed and recorded in Table 4, with their distributions visualized in Figure 2 of the main text.
-This experimental design allows us to analyze the inherent properties of different matrices and their roles in feature representation.
-C.2 Cross-Model Analysis of Matrix Properties
-We extend our analysis across three state-of-the-art text-to-image models: SDXL, SD3, and FLUX.1-dev, to validate the model-agnostic nature of our approach.
-As shown in Figure 14, Figure 16, and Figure 15, the layer-wise cosine similarity patterns remain consistent across different model architectures, demonstrating the generalizability of our QR-LoRA framework.
-Specifically:
-  • $Q$ Matrix Properties: Across all models, $Q$ matrices maintain remarkably high similarities (0.90-0.99), indicating their role as stable orthogonal bases for feature representation.
-  • $R$ Matrix Characteristics: The directly fine-tuned $R$ matrices show moderate but significant similarities (0.79-0.89), suggesting their inherited structural information from the original weights.
-  • $\Delta R$ Matrix Behavior: In contrast, $\Delta R$ matrices consistently exhibit substantially lower similarities (-0.17-0.25), validating their effectiveness in capturing task-specific features independently.
-Figure 11: User study interface. Screenshot of our evaluation interface where participants rate generated images on a 0-5 scale based on content preservation and style transfer quality.
-C.3 Statistical Analysis of Matrix Similarities
-The numerical results presented in Table 4 provide a detailed quantitative perspective on the similarity patterns. The data reveals several key insights:
-  • The maximum similarity values for $Q$ matrices consistently approach 1.0 (0.9998-0.9999), with minimum values remaining above 0.90, demonstrating the stability of the orthogonal basis.
-  • $R$ matrices show maximum similarities above 0.99 but with lower minimum values (0.79-0.89), reflecting their role in capturing task-relevant transformations while maintaining some degree of shared structure.
-  • $\Delta R$ matrices exhibit significantly lower similarity values, with maximums below 0.26 and minimums reaching -0.17, quantitatively confirming their capacity for independent feature representation.
-  • In comparison, traditional LoRA’s $A$ matrices show consistently high similarities (0.77-1.0), while $B$ matrices display moderate to low similarities (0.15-0.38), highlighting the advantages of our orthogonal decomposition approach.
-This comprehensive analysis validates the theoretical foundations of our QR-LoRA framework and demonstrates its effectiveness in achieving feature disentanglement across different model architectures.
-The consistent patterns observed in both layer-wise visualizations and statistical distributions support our design choice of focusing on $\Delta R$ updates for efficient and clean feature separation.
-Figure 12: Additional visualization results on different backbone models.
-Extended demonstration of QR-LoRA’s generation results across SDXL, SD3, and FLUX.1-dev models, further validating the consistent high-quality performance and model-agnostic nature of our approach.
-These supplementary examples provide a more comprehensive view of our method’s capabilities across diverse model architectures, complementing the results presented in Figure 6.
-Figure 13: Detailed experimental setup for matrix similarity comparison in Figure 2.
-We conduct experiments on 27 randomly selected image pairs to analyze the properties of different matrices in our QR-LoRA framework.
-For each image in the pairs, we train two types of LoRA adaptations: one using QR decomposition where both $Q$ and $R$ matrices are directly fine-tuned and another using the $\Delta R$ formulation.
-We then compute the cosine similarities between $Q$ matrices, $R$ matrices, and $\Delta R$ matrices across all injection layers in the SDXL model.
-Table 4: Numerical results of matrix similarity comparison in Figure 2.
-This table presents the detailed cosine similarity values for $Q$ matrices, $R$ matrices, $\Delta R$ matrices, $A$ matrices, and $B$ matrices from 27 pairs of image training data.
-For each image pair (i.e., sample index), we compute the cosine similarities across all injection layers in the SDXL model and record both the maximum ($Q_{max}$, $R_{max}$, $\Delta R_{max}$, $A_{max}$, $B_{max}$) and minimum ($Q_{min}$, $R_{min}$, $\Delta R_{min}$, $A_{min}$, $B_{min}$) similarity values.
-Sample Index $Q_{max}$ $Q_{min}$ $R_{max}$ $R_{min}$ $\Delta R_{max}$ $\Delta R_{min}$ $A_{max}$ $A_{min}$ $B_{max}$ $B_{min}$
-0 0.9999 0.9787 0.9997 0.8409 0.1657 -0.1139 1.0000 0.8046 0.3615 -0.0355
-1 0.9999 0.9786 0.9997 0.8253 0.1632 -0.1223 1.0000 0.7970 0.3249 -0.0625
-2 0.9999 0.9682 0.9997 0.8391 0.1135 -0.1038 1.0000 0.8161 0.3198 -0.0826
-3 0.9998 0.9311 0.9996 0.8104 0.1959 -0.1335 1.0000 0.8075 0.2322 -0.0565
-4 0.9998 0.9313 0.9996 0.7993 0.1241 -0.0818 1.0000 0.8020 0.2295 -0.0626
-5 0.9998 0.9055 0.9997 0.8143 0.1529 -0.0872 1.0000 0.8144 0.2226 -0.0558
-6 0.9998 0.9679 0.9997 0.8235 0.1566 -0.0581 1.0000 0.7917 0.2113 -0.0807
-7 0.9999 0.9680 0.9997 0.8129 0.1421 -0.0699 1.0000 0.8126 0.2145 -0.0783
-8 0.9999 0.9761 0.9997 0.8187 0.1257 -0.0913 1.0000 0.8053 0.1898 -0.2033
-9 0.9999 0.9709 0.9998 0.8441 0.1749 -0.1036 1.0000 0.8272 0.3044 -0.1188
-10 0.9999 0.9741 0.9998 0.8245 0.1500 -0.0783 1.0000 0.8163 0.2070 -0.0793
-11 0.9999 0.9792 0.9998 0.8833 0.1399 -0.0611 1.0000 0.8379 0.1525 -0.1278
-12 0.9999 0.9765 0.9997 0.8485 0.1816 -0.1094 1.0000 0.8253 0.2499 -0.0961
-13 0.9999 0.9793 0.9998 0.8216 0.2493 -0.0463 1.0000 0.8164 0.1490 -0.1374
-14 0.9999 0.9854 0.9999 0.8894 0.2067 -0.0933 1.0000 0.8332 0.1817 -0.1316
-15 0.9998 0.9724 0.9996 0.8410 0.1756 -0.1686 1.0000 0.8416 0.3820 -0.0801
-16 0.9998 0.9752 0.9996 0.8186 0.1851 -0.1050 1.0000 0.8361 0.2329 -0.0636
-17 0.9998 0.9807 0.9997 0.8818 0.1262 -0.0695 1.0000 0.8413 0.2125 -0.1462
-18 0.9999 0.9055 0.9996 0.8340 0.2539 -0.0760 1.0000 0.8346 0.3847 -0.2045
-19 0.9999 0.9054 0.9997 0.8408 0.1209 -0.0420 1.0000 0.8398 0.2905 -0.0762
-20 0.9999 0.9053 0.9996 0.8339 0.1705 -0.0540 1.0000 0.8298 0.2734 -0.0548
-21 0.9998 0.9045 0.9995 0.8040 0.2042 -0.0466 1.0000 0.7727 0.1746 -0.0739
-22 0.9998 0.9044 0.9995 0.8095 0.1441 -0.0716 1.0000 0.7714 0.1531 -0.1031
-23 0.9998 0.9045 0.9995 0.7992 0.1473 -0.0588 1.0000 0.7853 0.1558 -0.0861
-24 0.9998 0.9055 0.9996 0.8272 0.1711 -0.0512 1.0000 0.8034 0.1969 -0.0806
-25 0.9998 0.9053 0.9996 0.8207 0.1686 -0.0851 1.0000 0.8147 0.2685 -0.0401
-26 0.9998 0.9053 0.9996 0.8134 0.1719 -0.0850 1.0000 0.8108 0.2691 -0.0488
-Figure 14: Extended analysis of cosine similarities under SDXL model, complementing Figure 3.
-We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the SDXL backbone model to validate the generalization of our findings.
-The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results.
-Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns.
-This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
-Figure 15: Extended analysis of cosine similarities under FLUX.1-dev model, complementing Figure 3.
-We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the FLUX.1-dev backbone model to validate the generalization of our findings.
-The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results.
-Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns.
-This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
-Figure 16: Extended analysis of cosine similarities under SD3 model, complementing Figure 3.
-We replicate the similarity analysis experiment (sample index from 0 to 4) from Figure 3 using the SD3 backbone model to validate the generalization of our findings.
-The left panel shows the layer-wise cosine similarity trends across different training strategies, consistent with our main results.
-Additionally, we provide histogram distributions of similarity values in the right panel, offering a statistical perspective on the similarity patterns.
-This comprehensive visualization further supports our theoretical analysis by demonstrating that the observed matrix properties are consistent across different model architectures.
-Appendix D Additional Experimental Results and Discussion
-D.1 Ablation Studies
-Scaling coefficient.
-To comprehensively evaluate the robustness and flexibility of our QR-LoRA framework, we conduct extensive ablation studies on scaling coefficients across three different backbone models: SDXL, SD3, and FLUX.1-dev.
-Our experiments systematically explore coefficient combinations (i.e., $\lambda_{c},\lambda_{s}$) ranging from 0.5 to 1.0 with intervals of 0.1, resulting in 36 distinct combinations for each model.
-This granular analysis reveals several key insights about our method’s behavior and capabilities.
-As demonstrated in Figures 17, 18, and 19, our approach maintains consistent generation quality across different coefficient settings while enabling fine-grained control over the content-style fusion process.
-The smooth transitions observed between different coefficient combinations indicate that our QR decomposition effectively disentangles content and style features, allowing for intuitive and stable interpolation between these attributes.
-Notably, this behavior remains consistent across all three model architectures, validating the model-agnostic nature of our approach.
-Furthermore, our experiments demonstrate that our method exhibits robust performance across a wide range of scaling coefficients after obtaining the disentangled deltaR matrices.
-This scaling exploration serves as a deeper investigation into our method’s capabilities, rather than a limitation.
-It’s worth noting that such scaling strategies are generally applicable to any LoRA-based methods, showcasing the versatility of our approach in achieving various levels of content preservation and style transfer effects.
-Rank.
-To comprehensively evaluate the robustness of our method, we conduct a sensitivity analysis across different rank settings.
-As shown in Figure 10, our method demonstrates consistent performance and stable generation quality across various rank values, indicating its inherent robustness to this hyperparameter.
-We adopt rank=64 for all experiments in the main paper, as it strikes an optimal balance between computational efficiency and generation quality.
-This choice aligns well with common practices in the literature [49, 8] and proves to be a reliable setting for most applications.
-Figure 17: Comprehensive ablation study on scaling coefficients for content-style fusion based on SDXL.
-We demonstrate the generation results under various combinations of content and style scaling coefficients.
-The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations.
-Please zoom in to view details.
-Figure 18: Comprehensive ablation study on scaling coefficients for content-style fusion based on SD3.
-We demonstrate the generation results under various combinations of content and style scaling coefficients.
-The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations.
-Please zoom in to view details.
-Figure 19: Comprehensive ablation study on scaling coefficients for content-style fusion based on FLUX.1-dev.
-The experiments validate the flexibility and robustness of our approach in achieving fine-grained control over content-style fusion through different coefficient combinations.
-Please zoom in to view details.
-D.2 User Study
-To evaluate the perceptual quality of our results, we conducted a comprehensive user study with 47 participants (40 males and 7 females).
-The participants represented a diverse demographic, with ages ranging from 22 to 45 years (mean age: 31.2).
-Among the participants, 15 were computer vision researchers, 12 were professional designers, and 20 were general users with basic knowledge of image editing.
-This diverse composition ensured a balanced evaluation from both technical and aesthetic perspectives.
-Each participant was asked to rate the quality of generated images from different methods on a scale of 0-5, where 0 indicates poor quality and 5 indicates excellent quality.
-To facilitate the evaluation process, we designed an intuitive rating interface (shown in Figure 11) that allows participants to easily assess and score the generated images.
-The evaluation criteria focused on both content preservation and style transfer aspects.
-The ratings reported in Table 1 represent the average scores given by participants for each method.
-The consistently higher ratings achieved by our method (4.07 for SDXL, 4.14 for SD3, and 3.96 for FLUX.1-dev) compared to baseline approaches demonstrate the superior perceptual quality across different backbone models.
-D.3 Additional Qualitative Results
-To comprehensively demonstrate the effectiveness and versatility of our QR-LoRA framework, we present extensive qualitative results across different backbone models and diverse application scenarios.
-First, we conduct a thorough comparative analysis through Figure 20, which showcases our method’s superior performance against state-of-the-art approaches across SDXL, SD3, and FLUX.1-dev models.
-The results consistently demonstrate QR-LoRA’s ability to achieve better content preservation while maintaining high-fidelity style transfer, regardless of the underlying model architecture.
-Figure 20: Extended qualitative comparison results.
-Comparison of QR-LoRA against state-of-the-art methods on SDXL, SD3, and FLUX.1-dev models, demonstrating consistent superior performance across different scenarios.
-Please zoom in to view details.
-To further validate the model-agnostic nature of our approach, we present a comprehensive set of generation results across different backbone architectures.
-Figures 21, 22, and 23 showcase diverse generation results on SDXL, demonstrating our method’s capability in handling various challenging scenarios while maintaining high visual quality and style fidelity.
-The results span a wide range of content types and artistic styles, from intricate architectural details to subtle texture variations.
-Figure 21: Diverse generation results based on SDXL.
-Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
-Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics.
-Please zoom in to view details.
-Figure 22: Diverse generation results on SDXL.
-Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
-Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics.
-Please zoom in to view details.
-Figure 23: Diverse generation results on SDXL.
-Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
-Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics.
-Please zoom in to view details.
-Extending our analysis to other backbone architectures, Figure 24 presents generation results on SD3, while Figure 25 shows results using FLUX.1-dev as the backbone.
-These examples demonstrate that QR-LoRA maintains its effectiveness across different model architectures, consistently producing high-quality results that effectively balance content preservation and style transfer.
-Figure 24: Diverse generation results on SD3. Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
-Figure 25: Diverse generation results on FLUX.1-dev.
-Additional examples showcasing QR-LoRA’s capability in handling various style-content combinations.
-Our method successfully preserves intricate content details while effectively transferring diverse artistic styles, from painterly effects to digital art aesthetics.
-Please zoom in to view details.
-Collectively, these results validate the robustness and generalization capability of our approach across diverse scenarios.
-The consistent high-quality performance across different content types (e.g., portraits, objects) and artistic styles (e.g., watercolor, digital art) demonstrates QR-LoRA’s effectiveness in handling complex geometric structures, fine-grained textures, and dramatic style variations.
-This comprehensive evaluation further supports the model-agnostic nature of our method and its potential for broad application in various image generation and style transfer tasks.
-D.4 Ability to Re-contextualize
-Our method not only excels at fusing content and style but also enables further customization through text prompts.
-While maintaining the high-quality integration of content and style elements, our approach supports flexible scene manipulation through additional textual instructions, as shown in Figure 26.
-For instance, given a content image of a dog and a specific artistic style, we can generate variations like ”wearing a hat” or ”riding a bicycle” while preserving the desired style characteristics.
-This re-contextualization capability is particularly valuable in practical artistic applications, as it allows artists to explore diverse creative possibilities while maintaining consistent stylistic elements across different scenarios.
-Figure 26: Demonstration of our method’s re-contextualization capability.
-Given a content image and a style reference, our approach can generate variations through additional text prompts while maintaining style consistency.
-The examples show how different textual instructions (e.g., ”wearing a hat”, ”riding a bicycle”) guide the generation process to create diverse scenarios while preserving the original style characteristics.
-D.5 The End
-We anticipate that our model-agnostic, low-rank efficient fine-tuning approach will provide a versatile solution for fusion tasks requiring robust disentanglement properties, such as the fusion of arbitrary style and content.
-This stands in contrast to the field’s ongoing trend of pursuing specific disentanglement properties within particular models.
+Acknowledgments
+This research is funded by the Chongqing Natural Science Foundation Innovation and Development Joint Fund (Changan Automobile) (Grant No. CSTB2024NSCQ-LZX0157), China.